OpenStudy (anonymous):

McGrath earned a profit of Rs. 300 by selling 100 kg of a mixture of A and B types of rice, at a total price of Rs. 1,100. What was the proportion of A and B types of rice in the mixture if the cost prices of A and B are Rs. 10 and Rs. 5 per kg respectively?

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

a+b=100

OpenStudy (anonymous):

so price =800

OpenStudy (dumbcow):

good 10A +5B = 800

OpenStudy (anonymous):

800/10+5

OpenStudy (anonymous):

a=60,b=40

OpenStudy (dumbcow):

solve just like last problem

OpenStudy (dumbcow):

yes correct so ratio is a/b = 6/4 = 3/2